OpenStudy (baldymcgee6):

A fuse has a length of 10 meters and burns at a rate of two centimetres per second when lit. What length of fuse (in cm) remains after t seconds and when will it totally burned out?

OpenStudy (baldymcgee6):

@amistre64, @Hero

hero (hero):

Hint: D = rt

OpenStudy (baldymcgee6):

can we use calculus?

hero (hero):

You should always try to solve a problem without calculus if it isn't necessary

OpenStudy (baldymcgee6):

would it not be something like 10-0.02t

OpenStudy (amistre64):

it burns at a rate of 4 cms; so im sure we can if we need to

OpenStudy (baldymcgee6):

2cm*s^-1

OpenStudy (amistre64):

for some reason i read it as lit at both ends lol

hero (hero):

D - rt = 0

OpenStudy (baldymcgee6):

right, so 10m - 0.02cm*s^-1*t

OpenStudy (baldymcgee6):

D=10m - 0.02cm*s^-1*t

OpenStudy (amistre64):

\[\frac{df}{dt}=-.02\] \[f(t)=-.02t+C\]initial condition f(0)=10

OpenStudy (baldymcgee6):

right, so when will it be totally burned out?